### v1

# Access Hoare Logic

Abstract.
Following Hoare’s seminal invention, later called Hoare logic, to reason about correctness of computer programs, we advocate a related but fundamentally different approach to reason about access security of computer programs such as access control. We define the formalism, which we denote access Hoare logic, and present examples which demonstrate its usefulness and fundamental difference to Hoare logic. We prove soundness and completeness of access Hoare logic, and provide a link between access Hoare logic and standard Hoare logic.
Access control, blockchain, Hoare triple, access Hoare logic, Bitcoin, smart contracts
copyright: acmcopyrightjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: XXjournalnumber: Xarticle: 111publicationmonth: 8ccs: Theory of computation Hoare logicccs: Theory of computation Logic and verificationccs: Software and its engineering Software verificationccs: Security and privacy Logic and verificationccs: Security and privacy Software security engineering
1. Introduction
The correctness of programs has been a concern early on in the development of computers. The seminal contribution by Tony Hoare (Hoare, 1969) allows to reason about programs instruction by instruction using pre- and postconditions. It allows to ascertain that if the precondition is satisfied before execution, and program execution terminates, then the postcondition will be satisfied.
This form of ascertaining the correctness of programs has been and still is very successful. It is at the heart of several tools that are used in academia and industry to specify and verify software systems, such as SPARK (Adacore, 2025b) and Dafny (Dafny Development Team, 2025).
However, there are security properties that do not naturally fall within the framework of Hoare logic. Access control is the property of restricting access to a resource. It is a well-established area of computer security that includes identification, authorization, authentication, access approval, and auditing (Gollmann, 2011), where access approval is the process of deciding whether to grant or reject an access request. Access control has a broad range of applications, ranging from large distributed systems to local systems such as electronic keys. More recently, it gained importance in the context of blockchain and cryptocurrencies: Smart contracts are often written to control access to digital assets, such as cryptocurrencies, tokens, or other forms of digital assets managed by blockchain systems, thereby combining several properties of general access control. We will use the term access security to speak about such properties in general.
For a program to be access secure, we need to guarantee that the final state (e.g., access to a hotel room via electronic keys) will only be granted if a given set of conditions has been met before execution (e.g., possession of valid access codes). That is, for access security we are interested in the reverse direction of Hoare logic, which we will denote access Hoare logic (aHl), where we will reason from postconditions to preconditions. While normal Hoare logic expresses a precondition to be sufficient for the postcondition to hold after executing a program, access Hoare logic requires the precondition to be necessary. More precisely, access Hoare logic expresses that if the program starting in state $s$ terminates in a state $s^{\prime}$ which fulfills the postcondition, then it is necessary that state $s$ fulfills the precondition.
Another form of reversing Hoare logic has been proposed in the literature (de Vries and Koutavas, 2011) and further developed to deal with incorrectness of programs (O’Hearn, 2019). That formalism expresses that for each final state $s^{\prime}$ satisfying the postcondition there is a state $s$ satisfying the precondition such that the program, when started in state $s$, can terminate in state $s^{\prime}$. This is a different perspective, which is incompatible with our approach, to the best of our knowledge.
In this brief note, we define access Hoare triples and access Hoare logic and use it to describe access security for examples. We show that access Hoare logic is sound and complete for a standard, general interpretation of access Hoare triples. We provide a link between Hoare logic and access Hoare logic in that weakest preconditions for usual Hoare logic correspond to strongest preconditions for access Hoare logic on states where the program terminates.
Alhabadi et al. (Alhabardi et al., 2022) have recently discovered that weakest preconditions for Hoare logic can be used to express access security for Bitcoin scripts. Our observation here expands on those insights and provides a reason why that is the case.
2. Examples
We demonstrate our point of view by considering three examples: The first involves electronic keys, the second owning and transferring Bitcoins, and the third granting access depending on the presence of a key in a list of keys.
2.1. Access Security for Electronic Keys
Our first example is a digital key system in a hotel that uses cards to open doors, which was specified and verified by Jackson (Jackson, 2012) using Alloy, and by Nipkow (Nipkow, 2006) using Isabelle/HOL. In this system, the battery-powered door locks are not connected to a network. Instead, their state is controlled by cards that are inserted into them, or, in wireless systems, moved close to them. A lock holds an internal electronic key $\mathtt{dk}$ which controls its state. A card holds two electronic keys, $\mathtt{ck1}$ for the first key and $\mathtt{ck2}$ for the second. When a new guest receives a card (for example, from the hotel reception), it will contain the key corresponding to a previous guest as $\mathtt{ck1}$, and a new fresh key as $\mathtt{ck2}$. An access control program for the door lock suitable for use in such an electronic hotel key system should give access to the room in an access-secure way, by setting a variable $\mathtt{acc}$ to true: When the guest uses their card for the first time, access will be granted if the previous guest’s key is stored as the door key ($\mathtt{dk}{}=\mathtt{ck1}{}$), in which case the door key is updated to the new key $\mathtt{ck2}$, preventing the previous guest from gaining access to the room thereafter. The next time the new guest swipes their card, access will be granted because the second key on their card, $\mathtt{ck2}$, now matches the door key ($\mathtt{dk}{}=\mathtt{ck2}{}$).
For this example, a preconditon for verifying access security is given by $\mathtt{ck1}=\mathtt{dk}\lor\mathtt{ck2}=\mathtt{dk}$, expressing that either of the first key or the second key is the door key. A postcondition is $\mathtt{acc}=\mathtt{true}$, expressing that access is granted.
A program to solve this task may look as follows:
P0:  if not (dk == ck1)
       then acc := (dk == ck2)
       else dk := ck2; acc := true
However, there is an ambiguity in the way this code is written, in that it can be parsed in different ways. The intended way is that both instructions after else would be executed in the else part, i.e. the program would be parsed as
P1:  if not (dk == ck1)
       then {acc := (dk == ck2)}
       else {dk := ck2; acc := true}
Another reading would give higher priority to ’;’, leading to
P2:  {if not (dk == ck1)
       then {acc := (dk == ck2)}
       else {dk := ck2}};
     acc := true
Both P1 and P2 are correct w.r.t. the given pre- and postcondition in standard Hoare logic. However, only P1 is access secure w.r.t. the given pre- and postcondition, while P2 is not: the given precondition is not necessary for the postcondition to hold after executing P2, as P2 sets acc to true in any case.
2.2. Access Security for Bitcoin
The cryptocurrency Bitcoin is managed by a blockchain that stores transactions of Bitcoins on a ledger. The transfer of Bitcoins from a current owner, Bob, to a recipient, Alice, via transactions is controlled by small programs called scripts: Bob’s Bitcoins are protected by a locking script, while the new transaction that tries to transfer the Bitcoins to Alice contains an unlocking script. To transfer the Bitcoins to Alice, the unlocking script, followed by the locking script, are executed; if this execution succeeds (by producing the output true), the transaction is successful and the Bitcoins are transferred to Alice; otherwise, the transaction fails and the Bitcoins will not be transferred.
In this context, the focus is on the access security of the locking script: In any successful run of the unlocking and then the locking script, the unlocking script is required to finish in a state which fulfills a suitable precondition, see Fig. 1. For a typical locking script like the standard locking script $\mathtt{scriptPubKey}_{\mathtt{pubKeyHash}}$ (Bitcoin Community, 2025), a typical precondition would express that an address and a signature are provided, where the address hashes to a value stored in the locking script, and the signature matches the address.
direction of executionunlocking scriptlocking scriptpreconditionpostcondition
Figure 1. Executing the combined unlocking and locking script
2.3. Access Security for While Loops
We consider a program to grant access by setting a variable acc to true if a given passkey p is amongst a list of stored passkeys L, where lh(L) denotes the length of list L, and el(i,L) denotes the i-th element of list L. The output of this program (i.e. the variable $\mathtt{acc}$) may then be used by another system to e.g. open a door.
  inputs:  p : Keys
           L : List(Keys)
  outputs: acc : Boolean
  program CheckList
    i := 1;
    acc := false;
    while i<=lh(L) do
      if el(i,L) == p then acc := true endif;
      i := i+1;
    endwhile
Suitable pre- and postconditions for certifying the access security of this program should reflect the requirement that access should only be granted if a given passkey $\mathtt{p}$ is included in a list of stored passkeys $\mathtt{L}$. A suitable postcondition is given by $\mathtt{acc}=\mathtt{true}$. A suitable precondition for when to give access is that $\mathtt{p}$ occurs in $\mathtt{L}$, which may be expressed as $\exists j\leq\mathtt{lh}(\mathtt{L})(\mathtt{el}(j,\mathtt{L})=\mathtt{p})$. A suitable invariant for the while loop that reflects the intended behaviour of the program may be $\mathtt{acc}=\mathtt{true}\ \lor\ \exists j\leq\mathtt{lh}(\mathtt{L})(j\geq i\land\mathtt{el}(j,\mathtt{L})=\mathtt{p})$ where $i$ denotes the current index of the while loop. The above loop invariant reflects the intention that within the while loop, to reach the postcondition it is necessary that either the postcondition has been reached already ($\mathtt{acc}=\mathtt{true}$), or that a matching passkey in the list $\mathtt{L}$ will be found in the remaining steps of the loop ($\exists j\leq\mathtt{lh}(\mathtt{L})(j\geq i\land\mathtt{el}(j,\mathtt{L})=\mathtt{p})$).
3. Access Hoare Triples
Hoare logic is using Hoare triples to describe how the execution of code impacts on the state of the machine. The Hoare triple $\{P\}\,C\,\{Q\}$ consists of assertions $P$ and $Q$, and a piece of code $C$. $P$ is called the precondition, $Q$ the postcondition. $\{P\}\,C\,\{Q\}$ expresses that when $P$ is true and $C$ is executed and terminates then $Q$ is true. In other words, $P$ is a sufficient condition that $Q$ holds after successfully executing $C$. It can be expressed more formally by quantifying over all possible states:
(1) $\forall s,s^{\prime}[s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\land P(s)\rightarrow Q(s^{\prime})]$
where a state is a mapping of variables to values, and $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$ expresses that the execution of $C$ takes the executing machine from state $s$ to state $s^{\prime}$.
We are interested in the opposite direction, that $P$ is a necessary condition for $Q$ to hold after executing $C$. We again use a triple $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare triple, consisting again of a precondition $P$, a postcondition $Q$, and a program $C$. However, now $\langle P\rangle\,C\,\langle Q\rangle$ expresses that when the execution of $C$ results in $Q$ being true, then $P$ must have been true before, which can be expressed more formally as
(2) $\forall s,s^{\prime}[s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\wedge Q(s^{\prime})\rightarrow P(s)]$
In this case, we also say that $\langle P\rangle\,C\,\langle Q\rangle$ is true or valid.
There is a close relationship between access Hoare triples and Hoare triples by introducing negation:
Theorem 3.1.
$\langle P\rangle\,C\,\langle Q\rangle$ is equivalent to $\{\neg P\}\,C\,\{\neg Q\}$
Proof.
This follows since $Q(s^{\prime})\rightarrow P(s)$ is equivalent to $\neg P(s)\rightarrow\neg Q(s^{\prime})$. ∎
At this point we could introduce a calculus for access Hoare logic indirectly as negated Hoare logic in the sense of Theorem 3.1. However, it is of advantage to have a direct calculus for access Hoare logic. For example, a calculus for access Hoare logic would form the basis for systems that can verify access security in the sense of access Hoare logic using tools like theorem provers, similar to SPARK (Adacore, 2025b) and Dafny (Dafny Development Team, 2025) for verifying correctness based on standard Hoare logic. A reduction from access Hoare logic to standard Hoare logic, as outlined above, would introduce the contraposition of implications in verification proofs, i.e. $\neg B\rightarrow\neg A$ instead of $A\rightarrow B$. [^1] In more detail, a proof of an access Hoare triple $\langle P\rangle\,C\,\langle Q\rangle$ will usually involve subproofs of access Hoare triples of the form $\langle P^{\prime}\rangle\,C^{\prime}\,\langle Q^{\prime}\rangle$ and $\langle P^{\prime\prime}\rangle\,C^{\prime\prime}\,\langle Q^{\prime\prime}\rangle$, for certain subprograms $C^{\prime}$ and $C^{\prime\prime}$, where $C^{\prime}$ is immediately followed by $C^{\prime\prime}$. In such a situation, arguing access Hoare logic style from post- to preconditions requires proving the implication $P^{\prime\prime}\rightarrow Q^{\prime}$ from the precondition $P^{\prime\prime}$ of the second subproof to the postcondition $Q^{\prime}$ of the first. If we consider this situation by transformation to Hoare logic, we are faced with $\{\neg P\}\,C\,\{\neg Q\}$ and subproofs of $\{\neg P^{\prime}\}\,C^{\prime}\,\{\neg Q^{\prime}\}$ and $\{\neg P^{\prime\prime}\}\,C^{\prime\prime}\,\{\neg Q^{\prime\prime}\}$, now having to prove the implication $\neg Q^{\prime}\rightarrow\neg P^{\prime\prime}$ while arguing Hoare logic style from pre- to postconditions. For example, a straightforward implication such as $x=2\rightarrow x+2=4$ may translate into its contrapositive $\neg(x+2=4)\rightarrow\neg(x=2)$. While logically equivalent, [^2] In classical logic the latter is more difficult to understand and deal with in verification tasks. In general, this may even break verifiability. [^3] SPARK (Adacore, 2025b), which supports Hoare logic, tries to prove formulas generated by a verification process using an SMT solver. If that fails, SPARK offers to use the interactive theorem prover Coq/Rocq (Adacore, 2025a) instead. Another option when facing unprovability in SPARK, is to analyze formulas manually in Why3 (Why3, 2025; Adacore, 2025c), in order to find a reason for unprovability. In the last two cases, the syntactical form of formulas matters: For example, Coq/Rocq is based on intuitionistic logic, hence transformation of formulas need to respect intuitionistic logic.
We will introduce a direct calculus for access Hoare logic by stating its rules similar to those for Hoare logic. Soundness and completeness of such a calculus could be proven using soundness and completeness of Hoare logic and Theorem 3.1. As it turns out, proving soundness and completeness directly does not take much longer, and has the added benefit of motivating the rules, and keeping the paper selfcontained. Furthermore, it also supports a more direct and intuitive understanding of the rules of access Hoare logic as a formal method for access security.
Remark 3.2.
Following the previous discussion, we analyze Theorem 3.1 from the perspective of intuitionistic logic. The direction $\langle P\rangle\,C\,\langle Q\rangle\rightarrow\{\neg P\}\,C\,\{\neg Q\}$ is valid in intuitionistic logic, while for the reverse direction we only obtain $\{\neg P\}\,C\,\{\neg Q\}\to\langle\neg\neg P\rangle\,C\,\langle\neg\neg Q\rangle$. Observe that $\neg\neg P\rightarrow P$ is valid in classical logic, but not in intuitionistic logic in general.
4. The Calculus
We define the calculus for $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare logic, following that for Hoare logic (Hoare, 1969), and show that it is sound with respect to the interpretation of access Hoare triples as given in (2) in the previous section. The development will focus on the $\mathtt{while}$ language similar to (Hoare, 1969), a standard imperative language. We will show completeness in Section 6.
4.1. Axiom for empty statement
The empty statement $\mathtt{skip}$ does not change the state of the program, thus whatever holds after $\mathtt{skip}$ must have held before.
        $\langle P\rangle\,\mathtt{skip}\,\langle P\rangle$
Soundness of this rule is immediate.
4.2. Axiom scheme of assignment
An assignment of the from
$V:=E$
with $V$ a variable and $E$ an expression without side effects, but possibly containing $V$, updates variable $V$ with the result of evaluating $E$. Any assertion $P$ that is true after the assignment is made must also have been true of the value resulting from the evaluation of $E$.
          $\langle P[E/V]\rangle\,V:=E\,\langle P\rangle$
Here $P[E/V]$ denotes the result of replacing any free occurrence of $V$ in $P$ by $E$.
Soundness follows from the same considerations as for standard Hoare logic: Let $s,s^{\prime}$ be such that $s\stackrel{{\scriptstyle V:=E}}{{\longmapsto}}s^{\prime}$, then $P[E/V](s)$ is equivalent to $P(s^{\prime})$.
4.3. Consequence rule
The consequence rule is the main deviation from Hoare’s original calculus, reflecting that we are interested in expressing that the precondition is necessary for the postcondition. Thus, contrary to standard Hoare logic, we allow to weaken the precondition, and to strengthen the postcondition.
$P_{1}\leftarrow P_{2}$     $\langle P_{2}\rangle\,S\,\langle Q_{2}\rangle$     $Q_{2}\leftarrow Q_{1}$          $\langle P_{1}\rangle\,S\,\langle Q_{1}\rangle$
For soundness, let $s,s^{\prime}$ be such that $s\stackrel{{\scriptstyle S}}{{\longmapsto}}s^{\prime}$ and $Q_{1}(s^{\prime})$ hold. We need to show that $P_{1}(s)$ is true, assuming validity of the premises of this rule. From $Q_{1}(s^{\prime})$ we obtain $Q_{2}(s^{\prime})$ from the right premise. Validity of the middle premise then shows $P_{2}(s)$, from which the validity of the left premise implies $P_{1}(s)$.
4.4. Rule of composition
As in standard Hoare logic, $S;T$ denotes the sequential composition of $S$ and $T$, where $S$ executes prior to $T$.
$\langle P\rangle\,S\,\langle R\rangle$     $\langle R\rangle\,T\,\langle Q\rangle$        $\langle P\rangle\,S;T\,\langle Q\rangle$
For soundness, let $s,s^{\prime}$ be such that $s\stackrel{{\scriptstyle S;T}}{{\longmapsto}}s^{\prime}$ and $Q(s^{\prime})$ hold. We need to show that $P(s)$ is true, assuming validity of the premises of this rule. Let $s^{\prime\prime}$ such that $s\stackrel{{\scriptstyle S}}{{\longmapsto}}s^{\prime\prime}$ and $s^{\prime\prime}\stackrel{{\scriptstyle T}}{{\longmapsto}}s^{\prime}$. Validity of the right premise yields $R(s^{\prime\prime})$, from which validity of the left shows $P(s)$.
4.5. Conditional rule
The conditional rule also differs from standard Hoare logic. The assertion $B$ is used to ‘choose a branch’ corresponding to the ‘then’ or ‘else’ part of $\mathtt{if}$ B $\mathtt{then}$ S $\mathtt{else}$ T. For example, the left premise of the rule, in both standard and access Hoare logic, expresses ‘if $B$ then $(P)S(Q)$’: For standard Hoare logic, this expands to ’if $B$ then if $P$ and $S$ execute successfully, then $Q$’, which is logically equivalent to ’if $B\land P$ and $S$ execute successfully, then $Q$’. For access Hoare logic, the expression expands to ’if $B$ then if $Q$ and $S$ execute successfully, then $P$’, which is logically equivalent to ’if $Q$ and $S$ executes successfully, then $B\rightarrow P$’. Using a similar consideration for the right premise, the conditional rule takes the following form:
$\langle B\rightarrow P\rangle\,S\,\langle Q\rangle$     $\langle\neg B\rightarrow P\rangle\,T\,\langle Q\rangle$      $\langle P\rangle\,\mathtt{if}\ B\ \mathtt{then}\ S\ \mathtt{else}\ T\,\langle Q\rangle$
To formally argue for soundness of this rule, let $s,s^{\prime}$ be such that $s\stackrel{{\scriptstyle\mathtt{if}\ B\ \mathtt{then}\ S\ \mathtt{else}\ T}}{{\longmapsto}}s^{\prime}$ and $Q(s^{\prime})$ hold. We need to show that $P(s)$ holds, assuming validity of the two premises.
If $B(s)$ is true, then $s\stackrel{{\scriptstyle S}}{{\longmapsto}}s^{\prime}$, so $B(s)\rightarrow P(s)$ is true by validity of the left premise. Thus, $P(s)$ holds. Otherwise, $B(s)$ is false, and $s\stackrel{{\scriptstyle T}}{{\longmapsto}}s^{\prime}$, hence $\neg B(s)\rightarrow P(s)$ is true by validity of the right premise. Again, $P(s)$ holds.
4.6. While rule
Similarly to the conditional rule, the precondition of the premise changes from conjunction to implication compared to standard Hoare logic. Furthermore, we weaken the postcondition to only condition reachable states, namely on those satisfying $\neg B$ – this is needed to be able to prove completeness in Section 6.
     $\langle B\rightarrow P\rangle\,S\,\langle P\rangle$   $\langle P\rangle\,\mathtt{while}\ B\ \mathtt{do}\ S\,\langle\neg B\rightarrow P\rangle$
To prove soundness of this rule, let $s,s^{\prime}$ be such that $s\stackrel{{\scriptstyle\mathtt{while}\ B\ \mathtt{do}\ S}}{{\longmapsto}}s^{\prime}$ and $\neg B(s^{\prime})\rightarrow P(s^{\prime})$ hold. We need to show that $P(s)$ holds, assuming validity of the premise.
As $\mathtt{while}$ B $\mathtt{do}$ S terminates, there are some $k$ and states $s=s_{0}$, …, $s_{k}=s^{\prime}$ such that
$\forall i<k[B(s_{i})\land s_{i}\stackrel{{\scriptstyle S}}{{\longmapsto}}s_{i+1}]$
Furthermore, termination implies $\neg B(s_{k})$, hence $P(s_{k})$ by assumption. Hence, $s_{k-1}\stackrel{{\scriptstyle S}}{{\longmapsto}}s_{k}$ and validity of the premise $\langle B\rightarrow P\rangle\,S\,\langle P\rangle$ show $B(s_{k-1})\rightarrow P(s_{k-1})$. As we also have $B(s_{k-1})$, we obtain $P(s_{k-1})$ by modus ponens. Inductively, we obtain $P(s_{0})$ as required.
5. Weakest and Strongest Preconditions
The aim of this section is to show a connection between Hoare triples and access Hoare triples, namely that modulo termination, weakest preconditions for Hoare triples, and strongest preconditions for access Hoare triples coincide. We start by repeating standard definitions of weakest preconditions for Hoare triples, and provide a definition of strongest preconditions for access Hoare triples. We then give explicit characterizations of weakest preconditions for Hoare triples and strongest preconditions for access Hoare triples. We conclude by showing that, for a fixed program and post condition, the strongest precondition for Hoare triples equals to the intersection of the weakest precondition for Hoare Triples with the set of states on which the program terminates.
Definition 5.1 (Weakest precondition for Hoare triples).
Given a program $C$ and a postcondition $Q$, a weakest precondition is a predicate $P^{\prime}$ such that
$\forall P,\,\{P\}\,C\,\{Q\}\ \Leftrightarrow\ P\rightarrow P^{\prime}$
Definition 5.2 (Strongest precondition for access Hoare triples).
Given a program $C$ and a postcondition $Q$, a strongest precondition is a predicate $P^{\prime}$ such that
$\forall P,\,\langle P\rangle\,C\,\langle Q\rangle\ \Leftrightarrow\ P^{\prime}\rightarrow P$
It follows immediately from the definitions that weakest and strongest preconditions are preconditions for Hoare logic and access Hoare logic, respectively, and therefore are unique.
Lemma 5.3.
Given a program $C$ and a postcondition $Q$, let $P^{w}$ be defined by
$P^{w}(s)\ \Leftrightarrow\ \forall s^{\prime},\,s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\rightarrow Q(s^{\prime})$
Then $P^{w}$ is the weakest precondition for $C$ and $Q$ w.r.t. Hoare triples.
Proof.
To show that $P^{w}$ is the weakest precondition, it we need to show that
$\forall P,\,\{P\}\,C\,\{Q\}\Leftrightarrow P\rightarrow P^{w}$
Let $P$ be given. For the direction from right to left, assume $P\rightarrow P^{w}$. Let $s,s^{\prime}$ be given with $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$ and $P(s)$, hence $P^{w}(s)$. By the definition of $P^{w}$ and the choice of $s^{\prime}$, we have $Q(s^{\prime})$. This shows $\{P\}\,C\,\{Q\}$.
For the other direction, assume $\{P\}\,C\,\{Q\}$ and let $s$ with $P(s)$ be given. We need to show $P^{w}(s)$. To this end, assume $s^{\prime}$ such that $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$. By $\{P\}\,C\,\{Q\}$ we obtain $Q(s^{\prime})$, as required. Together, this shows $P\rightarrow P^{w}$. ∎
Lemma 5.4.
Given a program $C$ and a postcondition $Q$, let $P^{s}$ be defined by
$P^{s}(s)\ \Leftrightarrow\ \exists s^{\prime},\,s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\land Q(s^{\prime})$
Then $P^{s}$ is the strongest precondition for $C$ and $Q$ w.r.t. access Hoare triples.
Proof.
To show that $P^{s}$ is the strongest precondition, we need to show that
$\forall P,\,\langle P\rangle\,C\,\langle Q\rangle\Leftrightarrow P^{s}\rightarrow P$
Let $P$ be given. For the direction from right to left, assume $P^{s}\rightarrow P$. Let $s,s^{\prime}$ be given with $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$ and $Q(s^{\prime})$. Then $P^{s}(s)$ by the definition of $P^{s}$. Hence $P(s)$ by assumption. This shows $\langle P\rangle\,C\,\langle Q\rangle$.
For the other direction, assume $\langle P\rangle\,C\,\langle Q\rangle$ and let $s$ with $P^{s}(s)$ be given. By the definition of $P^{s}$, there exists $s^{\prime}$ such that $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$ and $Q(s^{\prime})$. By $\langle P\rangle\,C\,\langle Q\rangle$ we obtain $P(s)$. This shows $P^{s}\rightarrow P$. ∎
Corollary 5.5.
Let $P^{w}$ be the weakest precondition for $C$ and $Q$ w.r.t. Hoare triples, and $P^{s}$ be the strongest precondition for $C$ and $Q$ w.r.t. access Hoare triples. Then $P^{s}$ is equal to the intersection of $P^{w}$ with the set of states in which $C$ terminates.
Proof.
Let $T$ be the set of states in which $C$ terminates. Then we can argue
$\begin{array}[t]{@{}lcl}P^{s}(s)&\Leftrightarrow&\exists s^{\prime},\,s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\land Q(s^{\prime})\\ &\Leftrightarrow&(\exists s^{\prime},\,s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime})\land(\forall s^{\prime},\,s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\to Q(s^{\prime}))\\ &\Leftrightarrow&T(s)\land P^{w}(s)\end{array}$
where we have used that $C$ is deterministic. ∎
Corollary 5.6.
For programs that terminate on all inputs, weakest preconditions for Hoare logic and strongest preconditions for access Hoare logic coincide.
Remark 5.7.
  1. (a)
    The assertions of Lemmas 5.3, 5.4, and Corollaries 5.5, 5.6 hold in a general, abstract setting as long as the execution relation is deterministic, that is, they hold for any set $S$ (representing states) and any relation $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$ that satisfies
    $\forall s,s^{\prime},s^{\prime\prime},\ s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\land s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime\prime}\to s^{\prime}=s^{\prime\prime}\kern 5.0pt.$
  2. (b)
    Corollary 5.5 characterizes $P^{s}$. We can also characterize $P^{w}$, using a similar proof: $P^{w}$ consists of all states $s$ which satisfy that if $C$ terminates on $s$, then $s$ is in $P^{s}$.
6. Soundness and Completeness
Access Hoare logic, denoted aHl, is given by access Hoare triples and their semantics as defined in Section 3, and by the set of rules defined in Section 4. When introducing the rules, we showed already that all axioms and rules of aHl are sound. This implies the soundness of provable access Hoare triples:
Theorem 6.1 (Soundness of aHl).
If $\langle P\rangle\,C\,\langle Q\rangle$ is provable in aHl, then $\langle P\rangle\,C\,\langle Q\rangle$ is true.
We will now show the converse, that aHl is also complete:
Theorem 6.2 (Completeness of aHl).
If $\langle P\rangle\,C\,\langle Q\rangle$ is true, then $\langle P\rangle\,C\,\langle Q\rangle$ is provable in aHl.
We could prove this theorem by relying on the completeness of Hoare logic and Theorem 3.1. Since, as discussed before, we think there is value in treating access Hoare logic directly, we give here a direct proof.
To this end, we prove a slightly stronger proposition. In the previous section we have defined the strongest precondition for program $C$ and postcondition $Q$, here we define $P^{s}_{C}(Q)$ as an operator on $Q$ that is indexed by $C$:
$P^{s}_{C}(Q)\ :=\ \{s\colon\exists s^{\prime},\,s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\wedge Q(s^{\prime})\}$
Proposition 6.3.
For any program $C$ and predicate $Q$, $\langle P^{s}_{C}(Q)\rangle\,C\,\langle Q\rangle$ is provable in aHl.
This proposition implies completeness: Assume $\langle P\rangle\,C\,\langle Q\rangle$ is true. As $P^{s}_{C}(Q)$ is the strongest precondition for $C$ and $Q$, we obtain
$P^{s}_{C}(Q)\subseteq P$
The previous proposition shows that $\langle P^{s}_{C}(Q)\rangle\,C\,\langle Q\rangle$ is provable in aHl. Thus, using the consequence rule, we obtain that $\langle P\rangle\,C\,\langle Q\rangle$ is provable in aHl.
Proof of Proposition 6.3.
We prove the assertion by structural induction on $C$. For the empty statement $\mathtt{skip}$ we compute that $P^{s}_{\mathtt{skip}}$ is the identity function on predicates. Hence, the assertion follows immediately from the axiom for the empty statement.
In the case of assignment $V:=E$, we compute
$s\in P^{s}_{V:=E}(Q)\ \$$\Leftrightarrow\ \ \exists s^{\prime},\,s\stackrel{{\scriptstyle V:=E}}{{\longmapsto}}s^{\prime}\wedge Q(s^{\prime})$
$\Leftrightarrow\ \ s\in Q[E/V]$
Hence, $P^{s}_{V:=E}(Q)\ =\ Q[E/V]$ and the assertion again follows immediately from the corresponding axiom.
In case of sequential composition $S;T$, we can form the following aHl derivation by applying the rule of composition to the induction hypotheses for $S$ and $T$:
      i.h. $\langle P^{s}_{S}(P^{s}_{T}(Q))\rangle\,S\,\langle P^{s}_{T}(Q)\rangle$         i.h. $\langle P^{s}_{T}(Q)\rangle\,T\,\langle Q\rangle$          $\langle P^{s}_{S}(P^{s}_{T}(Q))\rangle\,S;T\,\langle Q\rangle$
An easy calculation shows $P^{s}_{S;T}=P^{s}_{S}\circ P^{s}_{T}$, hence the assertion follows.
In the case of the conditional statement $C=\mathtt{if}\ B\ \mathtt{then}\ S\ \mathtt{else}\ T$, let $P=P^{s}_{C}(Q)$. We claim
(3) $P^{s}_{S}(Q)$ $\rightarrow(B\rightarrow P)$
(4) $P^{s}_{T}(Q)$ $\rightarrow(\neg B\rightarrow P)$
For (3), assume $s\in P^{s}_{S}(Q)$ and $B(s)$. By definition, there exists $s^{\prime}$ such that $s\stackrel{{\scriptstyle S}}{{\longmapsto}}s^{\prime}\wedge Q(s^{\prime})$. As $B(s)$ holds, we obtain $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}\wedge Q(s^{\prime})$, thus $s\in P^{s}_{C}(Q)=P$. Similar for (4).
Using (3) and (4) and the induction hypotheses for $S$ and $T$, we can form the following derivation:
     (3) $P^{s}_{S}(Q)\rightarrow(B\rightarrow P)$         i.h. $\langle P^{s}_{S}(Q)\rangle\,S\,\langle Q\rangle$          $\langle B\rightarrow P\rangle\,S\,\langle Q\rangle$          (4) $P^{s}_{T}(Q)\rightarrow(\neg B\rightarrow P)$         i.h. $\langle P^{s}_{T}(Q)\rangle\,T\,\langle Q\rangle$          $\langle\neg B\rightarrow P\rangle\,T\,\langle Q\rangle$                      $\langle P\rangle\,\mathtt{if}\ B\ \mathtt{then}\ S\ \mathtt{else}\ T\,\langle Q\rangle$
For the while statement $C=\mathtt{while}\ B\ \mathtt{do}\ S$, let $P=P^{s}_{C}(Q)$. We claim
(5) $P^{s}_{S}(P)\rightarrow(B\rightarrow P)$
To prove (5), assume that $s\in P^{s}_{S}(P)$ and $B(s)$ hold. We need to show that $P(s)$ holds. From $s\in P^{s}_{S}(P)$ we obtain $s\stackrel{{\scriptstyle S}}{{\longmapsto}}s^{\prime}$ and $P(s^{\prime})$ for some $s^{\prime}$. $P(s^{\prime})$ implies $s^{\prime}\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime\prime}$ and $Q(s^{\prime\prime})$ for some $s^{\prime\prime}$. The first part, $s^{\prime}\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime\prime}$, implies that there are $s_{1},\dots,s_{k}$ such that $s_{1}=s^{\prime}$, $s_{k}=s^{\prime\prime}$, and
(6) $s_{i}\stackrel{{\scriptstyle S}}{{\longmapsto}}s_{i+1}\wedge B(s_{i})$
for $i=1,\dots,k{-}1$. Let $s_{0}=s$, then (6) also holds for $i=0$. Hence $s\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime\prime}$, which, together with $Q(s^{\prime\prime})$, shows $P(s)$. Thus, using the induction hypothesis, we can derive
    (5) $P^{s}_{S}(P)\rightarrow(B\rightarrow P)$         i.h. $\langle P^{s}_{S}(P)\rangle\,S\,\langle P\rangle$          $\langle B\rightarrow P\rangle\,S\,\langle P\rangle$        $\langle P\rangle\,\mathtt{while}\ B\ \mathtt{do}\ S\,\langle\neg B\rightarrow P\rangle$          $\star$ $Q\rightarrow(\neg B\rightarrow P)$               $\langle P\rangle\,\mathtt{while}\ B\ \mathtt{do}\ S\,\langle Q\rangle$
The final step is to show $\star$: Assuming $Q(s^{\prime})$ and $\neg B(s^{\prime})$, we need to show that $P(s^{\prime})$ holds. We observe that $\neg B(s^{\prime})$ implies $s^{\prime}\stackrel{{\scriptstyle C}}{{\longmapsto}}s^{\prime}$, which together with $Q(s^{\prime})$ shows $P(s^{\prime})$. ∎
We finish this section by drawing two interesting consequences that the following rules for conjunction and disjunction are admissible in access Hoare logic:
Conjunction rule
$\langle P_{1}\rangle\,C\,\langle Q_{1}\rangle$     $\langle P_{2}\rangle\,C\,\langle Q_{2}\rangle$      $\langle P_{1}\land P_{2}\rangle\,C\,\langle Q_{1}\land Q_{2}\rangle$
Disjunction rule
$\langle P_{1}\rangle\,C\,\langle Q_{1}\rangle$     $\langle P_{2}\rangle\,C\,\langle Q_{2}\rangle$      $\langle P_{1}\lor P_{2}\rangle\,C\,\langle Q_{1}\lor Q_{2}\rangle$
Similar rules are known to be admissible for standard Hoare logic (Gordon, 2022). The soundness of both rules for access Hoare logic follows immediately. Their admissibility is also true: If the premises for a rule are provable in aHl, then they are true by soundness. But then the conclusion is also true by the soundness of the rule, which is therefore provable in aHl by completeness.
7. Conclusion
In this brief note, we made the point that access security cannot be addressed naturally by standard Hoare logic. To overcome this issue, we introduced access Hoare logic where the reasoning is in the reverse direction compared to Hoare logic, namely from postconditions to preconditions. We gave three examples to demonstrate our point, one regarding electronic keys, a second considering the cryptocurrency Bitcoin, and a third about granting access involving a key and a list of keys. We introduced rules for access Hoare logic, and showed that these rules are sound and complete. We provided a link between access Hoare logic and Hoare logic by showing that weakest preconditions for Hoare logic coincides with strongest preconditions for access Hoare logic for total programs.
While access Hoare logic can be reduced to Hoare logic by introducing negation, we argued that this would have several negative consequences, including breaking verifiability using intuitionistic theorem provers — indeed, careful inspection shows that all proofs in this paper (except for Theorem 3.1, which needs to be replaced by Remark 3.2) only use intuitionistic logic, which we have confirmed using Agda (Agda Community, 2025) which is based on intuitionistic type theory. Details of that formalization will appear in a separate publication.
There are, of course, many areas that we did not touch upon. Basically, everything that has been investigated for Hoare logic should now be revisited in the context of access Hoare logic. Instead of aiming for a complete list, we mention two obvious areas. In our approach to soundness and completeness, we define the semantics of expressions and consequences in an informal way. There is a vast literature on ways to make this more precise, by considering formal semantics, their axiomatization and implications on Hoare logic (Apt, 1981). As a second area, the relationship to Dijkstra’s predicate transformer semantics and its variations need to be clarified (Dijkstra, 1975). Finally, detailed case studies are needed to demonstrate the usefulness of access Hoare logic for access security.
In connection to this paper, the authors have filed a patent entitled Verifying Access Security of a Computer Program (Beckmann and Setzer, 2025). This patent utilizes verification condition generation for access Hoare logic, which will be detailed in a separate publication.
Acknowledgements.
We would like to thank Fahad F. Alhabardi. This article builds upon and expands the use of weakest preconditions for access security, which were explored during his PhD research (Alhabardi, 2024). We would also like to thank John V Tucker for valuable comments on an earlier draft of this paper.